OpenStudy (anonymous):

More Financial Algebra help! Every six months, Jessa deposits $525 into an interest-bearing account to save for her children’s tuition. The interest rate on the account is 5.4% compounding semiannually. What is the present value of the investment if Jessa’s children leave for college in 11 years? $8,623.98 $6,665.39 $8,856.83 $7.025.32 I keep getting this as my answer: 1669.7549618492145787 which is nowhere close to any of the answer choices given above. ):

OpenStudy (anonymous):

This is how I have my problem set up FV = 525(1 + (5.4/100))^(2 * 11)

OpenStudy (anonymous):

@ganeshie8 Mind helping me again? lol :p

OpenStudy (anonymous):

Would I set it like this? FV = 525/(1 + (5.4/2))^(2 * 11)

OpenStudy (anonymous):

Never mind..that didn't work either :l

ganeshie8 (ganeshie8):

hint : its not a single deposit

OpenStudy (campbell_st):

well you need the formula for the future value of an annuity.... an annuity is a situation where the same amount is deposited under the same conditions.... the formula is \[FV = M \times \frac{(1 + r)^n -1}{r}\] M is the regular contribution... all the other information is the same as compound interest

ganeshie8 (ganeshie8):

every 6 months she is depositing, that means without considering interest itself, she should get : 525*2*11 money, your 1669 is very less

ganeshie8 (ganeshie8):

^^

OpenStudy (anonymous):

So, M = 525 r = 5.4 n = 2 Right?

OpenStudy (campbell_st):

oops I gave future value when you want the present value... so you need to take the future value answer... and divide it by (1 +r)^n to find out how much you would need to invest now to get the same amount in 11 years

OpenStudy (campbell_st):

the present value formula is \[PV = M \times \frac{(1 + r)^n -1}{r(1 + r)^n}\]

OpenStudy (campbell_st):

so if you calculate the future value of the annuity it will be $30902.90 so what amount do you need to invest now... for the same return uses the compound interest fromula \[30902.90 = P( 1 + 0.054)^{22}\] just make P the subject

OpenStudy (campbell_st):

well its a bit to much.... that will depend on then the $525 deposits are made.... at the beginning or end of each period....

ganeshie8 (ganeshie8):

when we're not explicitly told, we can assume its an ordinary annuity

OpenStudy (anonymous):

I am so lost...

OpenStudy (anonymous):

I tried both of your formulas and I still get the wrong answers

OpenStudy (anonymous):

...

OpenStudy (anonymous):

I'll just guess. Thanks anyways.

OpenStudy (campbell_st):

its the last answer

OpenStudy (anonymous):

No its not

OpenStudy (campbell_st):

well there you go... I just worked it out on a spread sheet and got the last answer...

OpenStudy (campbell_st):

ok... there in lies the problem.... the interest rate was 5.4% per annum and not 5.4% per half year.... now it makes more sense....

OpenStudy (anonymous):

So, the link I sent you was correct?

OpenStudy (campbell_st):

which is $8623.98

OpenStudy (anonymous):

I'll take that as a yes lol

OpenStudy (campbell_st):

so find the future value using the formula \[FV = 525 \times \frac{(1 + 0.027)^{22} -1}{0.027}\] which is $15497.40 so now use the compound interest formula with the futrue value \[15497.4 = P \times (1 + 0.027)^{22}\] solve for P and you get answer A

OpenStudy (anonymous):

Just checked and it was correct. (: Thanks

OpenStudy (campbell_st):

amazing how omitting p.a. changes the question

OpenStudy (anonymous):

Indeed lol